A neon atom (m=20.0u) makes a perfectly elastic collision with another atom at rest. After the impact, the neon atom travels away at a 18.4 ∘ angle from its original direction and the unknown atom travels away at a -69.2 ∘ angle.

What is the mass (in u) of the unknown atom? [Hint: You can use the law of sines.]
1
Expert's answer
2019-04-01T10:41:34-0400

From the conservation of momentum:


"m\vec{u}=m\vec{u'}+M\vec{v}"

Lets use the law of sines


"\frac{mu'}{\sin{69.2}}=\frac{Mv}{\sin{18.4}}=\frac{mu}{\sin{(180-69.2-18.4)}}"

"v=6.319uM^{-1}""u'=0.9356u"

From the conservation of energy:


"0.5mu^2=0.5mu'^2+0.5MV^2"

"0.5(20)u^2=0.5(20)(0.9356u)^2+0.5M(6.319uM^{-1})^2"

"M=16.0 u."